The mean of 12 observation is 39.7.If each observation is divided by 3 and 15 is added then find new mean

Grade:12th pass

2 Answers

Arun
4 years ago
Dear student
 
The changes which you make in every term, the same changes occur in mean.
 
Hence new mean = (39.7 /3) + 15
 = 13.233 + 15
 = 28.233..
Vikas TU
4 years ago
Dear student 
The above process is incorrect 
Mean = 39.7 
Sum = 39.7 *12 = 476.4 
476.4/3 + 15 = 173.8 
New mean = 173.8/12 =14.483
Good Luck